OpenStudy (anonymous):

Divide and simplify. frac{2y-10}{35}/{y-5}{49y}

OpenStudy (anonymous):

\[\frac {2y-10}{35}/\frac{y-5}{49y}\]

OpenStudy (anonymous):

\[\frac{2\times 49y}{35 }\]\[\frac{14y}{5}\]